OpenStudy (anonymous):

what is the limit as n approaches infinity of 1/n[(1/n)^9 + (2/n)^9 + (3/n)^9 + ... + (n/n)^9]

myininaya (myininaya):

how do you evaluate sum(i^9,i=1..n) do you know?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

please give a little more background. What class? Is this an infinite series?

OpenStudy (anonymous):

is there anything else to the question

OpenStudy (anonymous):

looks like a geometric series

OpenStudy (anonymous):

^yeah

OpenStudy (anonymous):

so find the ratio

OpenStudy (anonymous):

I have to think in terms of integrals to solve it.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

you need a power series representation and get the ratio from that

OpenStudy (anonymous):

the ratio is 2

OpenStudy (anonymous):

you want (1/n)^10 + (2/n)^10 + ...(

OpenStudy (anonymous):

i take that back

OpenStudy (anonymous):

the ratio of consecutive terms is not constant

OpenStudy (anonymous):

right

OpenStudy (anonymous):

this is an integral

OpenStudy (anonymous):

can yuo post the original question

OpenStudy (anonymous):

yeah thats what I asked

OpenStudy (anonymous):

oh its a riemann sum

OpenStudy (anonymous):

the limit is probably infinity

OpenStudy (anonymous):

its not a decreasing function

OpenStudy (anonymous):

no , if its a riemann sum its area under a curve

OpenStudy (anonymous):

that means it converges

OpenStudy (anonymous):

i just cant see the integral

OpenStudy (anonymous):

well do you see that n is finite there

OpenStudy (anonymous):

uphill, in that expression

OpenStudy (anonymous):

lim (n->oo) 1/n[(1/n)^9 + (2/n)^9 + (3/n)^9 + ... + (n/n)^9]

OpenStudy (anonymous):

n approaches infinity

OpenStudy (anonymous):

yes but each n is finite

OpenStudy (anonymous):

this is not a series

OpenStudy (anonymous):

this is a rieman sum

OpenStudy (anonymous):

maybe you need to set up an inequality

OpenStudy (anonymous):

S sub n + the limit of the integral of that is > the limit (of stuff)

OpenStudy (anonymous):

well the denominator is n^9

OpenStudy (anonymous):

no you dont need that

OpenStudy (anonymous):

lim (n->oo) 1/n[(1/n)^9 + (2/n)^9 + (3/n)^9 + ... + (n/n)^9] = 1/n [ (1^9 + 2^9 + 3^9 + ... n^9) / n^9 ]

OpenStudy (anonymous):

you need to find the sum formula for sum k^9 as for k = 1 to n

OpenStudy (anonymous):

let me try and make the problem look more neat and type it again

OpenStudy (anonymous):

give me medal

OpenStudy (anonymous):

i solve it

OpenStudy (anonymous):

when you pull out the /n^9 isn't it just a ratio of leading coefficients at that point? then you just have 1/n which goes to 0?

OpenStudy (anonymous):

leading coefficients?

OpenStudy (anonymous):

(L'pitals more or less)

OpenStudy (anonymous):

well i dont think you can apply Lhopital;s easily

OpenStudy (anonymous):

whichi is why u use leading coefficients

OpenStudy (anonymous):

we need to find a formula for 1^9 + 2^9 + ... n^9

OpenStudy (anonymous):

a closed form in terms of n

OpenStudy (anonymous):

okay. i know formulas for powers 1,2,3 but not 9

OpenStudy (anonymous):

and thanks you everyone for helping me out on this one. its pretty tough

OpenStudy (anonymous):

wait, can you state the original problem again

OpenStudy (anonymous):

sure. just a minute

myininaya (myininaya):

\[\int\limits_{0}^{1}x^{9}dx=x^{10}/10, x=0..1=1/10\]

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} (1/n)[(1/n)^9+(2/n)^9+(3/n)^9+...+(n/n)^9]\]

myininaya (myininaya):

deltax=[b-a]/n=[1-0]/n

OpenStudy (anonymous):

right that works

OpenStudy (anonymous):

or you can do the limit proof directly , find the sum of k^9

OpenStudy (anonymous):

thank you very much! i finally understand it. well more than i did beforehand anyways.